# Beyond Pairwise: Empowering LLM Alignment With Ranked Choice Modeling

### v1

@@ E.1.2 MNL-PO-Discrete
$\nabla_{\theta}\mathcal{L}_{\rm MNL-PO-Discrete}(\pi_{\theta})=$
@@
where the expectation is with respect to $(x,S,y_{w})\sim\mathcal{D}$. As MNL-PO-Discrete is a special case of MNL-PO-Top-k with ranking length $k=1$, we defer its derivation to the next section.
@@ E.1.3 MNL-PO-Top-k
$\nabla_{\theta}\mathcal{L}_{\rm MNL-PO-Top-k}(\pi_{\theta})=$
[…]
@@
$=-\beta\mathbb{E}\left[\sum_{i=1}^{k}\sigma\left({\rm log}\,\sum_{y_{j}\in S\setminus\{y_{1},\ldots,y_{i}\}}{\rm exp}(f_{\theta}(x,y_{j},y_{i}))\right)\cdot\sum_{y_{j}\in S\setminus\{y_{1},\ldots,y_{i}\}}\tfrac{\nabla_{\theta}\,{\rm log}\,\pi_{\theta}(y_{i}|x)-\nabla_{\theta}\,{\rm log}\,\pi_{\theta}(y_{j}|x)}{\sum_{y_{j^{\prime}}\in S\setminus\{y_{1},\ldots,y_{i}\}}{\rm exp}(f_{\theta}(x,y_{j^{\prime}},y_{i})-f_{\theta}(x,y_{j},y_{i}))}\right]$ (Eq. (19))
$=-\beta\mathbb{E}\left[\sum_{i=1}^{k}\sigma\left({\rm log}\,\sum_{y_{j}\in S\setminus\{y_{1},\ldots,y_{i}\}}{\rm exp}(f_{\theta}(x,y_{j},y_{i}))\right)\cdot\left[\nabla_{\theta}\,{\rm log}\,\pi_{\theta}(y_{i}|x)-\sum_{\begin{subarray}{c}y_{j}\in\\ S\setminus\{y_{1},\ldots,y_{i}\}\end{subarray}}\tfrac{\nabla_{\theta}\,{\rm log}\,\pi_{\theta}(y_{j}|x)}{\sum_{y_{j^{\prime}}\in S\setminus\{y_{1},\ldots,y_{i}\}}{\rm exp}(f_{\theta}(x,y_{j^{\prime}},y_{j}))}\right]\right]$
@@ E.2.3 Mallows-RMJ-PO-Top-k
$\mathcal{L}_{\rm Mallows-RMJ-PO-Top-k}(\pi_{\theta})=-\mathbb{E}_{(x,S,\mu^{k})\sim\mathcal{D}}\bigg[\log\phi(x)\big(\sum_{i=1}^{k-1}(|S|-i)\,\sigma(f_{\theta}(x,y_{i+1},y_{i}))\,+\sum_{y_{j}\in S\setminus\{y_{1},\ldots,y_{k}\}}\sigma(f_{\theta}(x,y_{j},y_{k}))\big)\bigg].$
@@
$=-\mathbb{E}\bigg[\log\phi(x)\big(\sum_{i=1}^{k-1}(|S|-i)\,\sigma^{\prime}(f_{\theta}(x,y_{i+1},y_{i}))\;\nabla_{\theta}f_{\theta}(x,y_{i+1},y_{i})\,+\sum_{y_{j}\in S\setminus\{y_{1},\ldots,y_{k}\}}\sigma^{\prime}(f_{\theta}(x,y_{j},y_{k}))\;\nabla_{\theta}f_{\theta}(x,y_{j},y_{k})\big)\bigg]$
@@ Appendix F Experimental Details
All experiments are implemented in Python 3.10.16 with PyTorch 2.6.0 on a server with 7 NVIDIA H100 GPUs each with 80 GB memory, equipped with Ubuntu 22.04.2 LTS.
@@
We provide descriptions of the training dataset and evaluation benchmarks in Tables 5 and 6, respectively. We construct a ranking-based preference set for ranked choice training in the following way. We generate five distinct responses for each prompt in the UltraFeedback dataset (Cui et al., 2023) using a sampling temperature of 0.8. We then score and rank these five responses with the Skywork-Reward-V2-Llama-3.1-8B reward model. In the pairwise setup, we adopt the literature’s tradition and use the top- and bottom- ranked responses as the assortment. For ranked choice training, we truncate each full ranking to generate the required data.
@@
Table 5: Overview of the UltraFeedback Dataset (Cui et al., 2023).
@@
Prompt Types
Open-ended dialogue
Instruction-following (writing, coding, summarization, translation)
Factual QA, adversarial/false QA
Standard NLP tasks (classification, reasoning, reading comprehension)
@@
AlpacaEval 2 805 GPT-4-Turbo GPT-4.1-mini Pairwise comparison LC & raw win rate
@@
F.1 Robustness Check
@@ Appendix G Qualitative Examples
  • Notes: In the Excel test, Mallows-RMJ-PO-Top-2 answered all three questions correctly, whereas DPO incorrectly calculated the mean in the last question using the average function.
@@
Table 11: Qualitative Example (Equation Solving)
Question: Solve this equation. $x^{3}-4x^{2}+6x-24=0$
Mallows-RMJ-PO-Pairwise DPO
A nice cubic equation! Let’s try to factor it: 1. Start with the given equation: $x^{3}-4x^{2}+6x-24=(x-3)(x^{2}-x-8)=0$ 2. Factor the quadratic inside: $x^{2}-x-8=(x-2)(x+4)=0$ 3. So we have: (x3)(x2)(x+4)=0(x-3)(x-2)(x+4)=0 4. This tells us that either: $x-3=0\;\Rightarrow\;x=3$ $x-2=0\;\Rightarrow\;x=2$ $x+4=0\;\Rightarrow\;x=-4$ Therefore, the solutions to the equation are $x=3,\quad x=2,\quad x=-4.$ A nice cubic equation! To solve this equation, we can try various methods. Here are a few approaches: 1. Method 1: Factoring
Let’s see if we can factor the left-hand side of the equation:
$x^{3}-4x^{2}+6x-24=(x-3)(x^{2}-x-8)=0$ Now, we need to factor the quadratic expression inside: $x^{2}-x-8=(x-2)(x+4)=0$ So, we have: (x3)(x2)(x+4)=0(x-3)(x-2)(x+4)=0 This tells us that: $x=3,\quad x=2,\quad x=-4$
2. Method 2: Numerical Methods or Formula
If factoring doesn’t work, we can use numerical methods or the cubic formula. For example:
$x=\tfrac{-b\pm\sqrt{b^{2}-3ac}}{3a}\;+\;$ $\tfrac{(2c\pm\sqrt{(b^{2}-3ac)(3a-2b)}+c^{3})}{3a^{2}}$ In this case, $a=1$, b=4b=-4, $c=6$. Plugging in the values gives messy expressions, and after simplification, we don’t get a neat closed form. The cubic formula often leads to complicated or approximate solutions.
Conclusion: The most straightforward approach is factoring, which yields the solutions: $x=3,\quad x=2,\quad x=-4$ These are the exact solutions to the equation: $x^{3}-4x^{2}+6x-24=0$
  • Notes: For the equation-solving task, Mallows-RMJ-PO-Pairwise provided the correct solution, whereas DPO gave an incorrect answer (method 2).

### v2

@@ E.1.2 MNL-PO-Discrete
$\nabla_{\theta}\mathcal{L}_{\rm MNL-PO-Discrete}(\pi_{\theta})$
@@
where the expectation is with respect to $(x,S,y_{w})\sim\mathcal{D}$. As MNL-PO-Discrete is a special case of MNL-PO-Topk with ranking length $k=1$, we defer its derivation to the next section.
@@ E.1.3 MNL-PO-Top-k
$\nabla_{\theta}\mathcal{L}_{\rm MNL-PO-Top-k}(\pi_{\theta})$
$=-\beta\mathbb{E}\bigg[\underbrace{\sum_{i=1}^{k}}_{\begin{subarray}{c}\text{ranked preference}\\ y_{1}\succ\ldots\succ y_{i}\mid S\ ;x\end{subarray}}\underbrace{\sigma\bigg({\rm log}\,\sum_{y_{j}\in S\setminus\{y_{1},\ldots,y_{i}\}}{\rm exp}(f_{\theta}(x,y_{j},y_{i}))\bigg)}_{\text{higher weight when reward deviates from preference}}$
$\quad\times\bigg[\nabla_{\theta}\,{\rm log}\,\pi_{\theta}(y_{i}|x)-\sum_{\begin{subarray}{c}y_{j}\in S\setminus\{y_{1},\ldots,y_{i}\}\end{subarray}}\tfrac{\nabla_{\theta}\,{\rm log}\,\pi_{\theta}(y_{j}|x)}{\underbrace{\sum_{\begin{subarray}{c}y_{j^{\prime}}\in S\setminus\{y_{1},\ldots,y_{i}\}\end{subarray}}{\rm exp}(f_{\theta}(x,y_{j^{\prime}},y_{j}))}_{\text{higher weight when reward is larger}}}\bigg]\bigg],$ (19)
@@
$=$ $-\beta\mathbb{E}\left[\sum_{i=1}^{k}\sigma\left({\rm log}\,\sum_{y_{j}\in S\setminus\{y_{1},\ldots,y_{i}\}}{\rm exp}(f_{\theta}(x,y_{j},y_{i}))\right)\right$.
$\quad\times\left.\sum_{y_{j}\in S\setminus\{y_{1},\ldots,y_{i}\}}\tfrac{\nabla_{\theta}\,{\rm log}\,\pi_{\theta}(y_{i}|x)-\nabla_{\theta}\,{\rm log}\,\pi_{\theta}(y_{j}|x)}{\sum_{y_{j^{\prime}}\in S\setminus\{y_{1},\ldots,y_{i}\}}{\rm exp}(f_{\theta}(x,y_{j^{\prime}},y_{i})-f_{\theta}(x,y_{j},y_{i}))}\right]$ (Eq. (20))
$=$ $-\beta\mathbb{E}\left[\sum_{i=1}^{k}\sigma\left({\rm log}\,\sum_{y_{j}\in S\setminus\{y_{1},\ldots,y_{i}\}}{\rm exp}(f_{\theta}(x,y_{j},y_{i}))\right)\right$.
$\quad\times\left.\bigg[\nabla_{\theta}\,{\rm log}\,\pi_{\theta}(y_{i}|x)-\sum_{\begin{subarray}{c}y_{j}\in S\setminus\{y_{1},\ldots,y_{i}\}\end{subarray}}\tfrac{\nabla_{\theta}\,{\rm log}\,\pi_{\theta}(y_{j}|x)}{\sum_{y_{j^{\prime}}\in S\setminus\{y_{1},\ldots,y_{i}\}}{\rm exp}(f_{\theta}(x,y_{j^{\prime}},y_{j}))}\bigg]\right]$
@@ E.2.2 Mallows-RMJ-PO-Discrete
$\Big[\underbrace{\nabla_{\theta}\log\pi_{\theta}(y_{i}|x)}_{\text{decrease likelihood of $y_{i}$}}-\underbrace{\nabla_{\theta}\log\pi_{\theta}(y_{w}|x)}_{\text{increase likelihood of $y_{w}$}}\Big]\bigg],$
@@ E.2.3 Mallows-RMJ-PO-Top-k
$\mathcal{L}_{\rm Mallows-RMJ-PO-Top-k}(\pi_{\theta})=$
$\quad-\mathbb{E}_{(x,S,\mu^{k})\sim\mathcal{D}}\bigg[\log\phi(x)\big(\sum_{i=1}^{k-1}(|S|-i)\,\sigma(f_{\theta}(x,y_{i+1},y_{i}))+\sum_{y_{j}\in S\setminus\{y_{1},\ldots,y_{k}\}}\sigma(f_{\theta}(x,y_{j},y_{k}))\big)\bigg].$
@@
$=-\mathbb{E}\bigg[\log\phi(x)\big(\sum_{i=1}^{k-1}(|S|-i)\,\sigma^{\prime}(f_{\theta}(x,y_{i+1},y_{i}))\;\nabla_{\theta}f_{\theta}(x,y_{i+1},y_{i})$
$\quad+\sum_{y_{j}\in S\setminus\{y_{1},\ldots,y_{k}\}}\sigma^{\prime}(f_{\theta}(x,y_{j},y_{k}))\;\nabla_{\theta}f_{\theta}(x,y_{j},y_{k})\big)\bigg]$
@@
$\quad\times\big(\underbrace{\nabla_{\theta}\log\pi_{\theta}(y_{i+1}\!\mid x)}_{\text{decrease likelihood of $y_{i+1}$}}-\underbrace{\nabla_{\theta}\log\pi_{\theta}(y_{i}\!\mid x)}_{\text{increase likelihood of $y_{i}$}}\big)$
@@ Appendix F Experimental Details
Our method is implemented by modifying the DPO Trainer and DPO Config in the TRL library. We include the modified codes and training scripts in the supplementary materials.
@@
All training runs were performed in Python 3.10.16 with PyTorch 2.6.0 on a server with 7 NVIDIA H100 GPUs each with 80 GB of memory, equipped with Ubuntu 22.04.2 LTS. The response-generation step in evaluations is performed on an NVIDIA Quadro RTX 6000 GPU with 24 GB of memory.
@@
We provide descriptions of the UltraFeedback train dataset in Table 6. We construct a ranking-based preference set for ranked choice training in the following way. We generate five distinct responses for each prompt in the UltraFeedback dataset (Cui et al., 2023) using a sampling temperature of 0.8. We then score and rank these five responses with the Skywork-Reward-V2-Llama-3.1-8B reward model. In the pairwise setup, we adopt the literature’s tradition and use the top- and bottom- ranked responses as the assortment. For ranked choice training, we truncate each full ranking to generate the required data.
The UltraFeedback test dataset contains 1961 prompts, and we use the chosen responses as the baseline. This dataset serves as an in-distribution test.
Benchmarks
We provide descriptions of the evaluation benchmarks in Table 7.
@@
Table 6: Overview of the UltraFeedback Train Dataset (Cui et al., 2023).
@@
Prompt Types Open-ended dialogue Instruction-following (writing, coding, summarization, translation) Factual QA, adversarial/false QA Standard NLP tasks (classification, reasoning, reading comprehension)
@@
AlpacaEval 2 805 GPT-4-Turbo GPT-4.1-mini Pairwise Comparison LC & Win Rate
@@
F.1 Training Time of MNL and Mallows-RMJ
Table 8: Training Time (Hours) of MNL and Mallows-RMJ.
Feedback MNL Mallows-RMJ
Pairwise Comparison 1.5 1.5
Discrete Choice 3.5 3.5
Top 2 Choice 3.5 3.5
F.2 Robustness Check
@@
F.3 More Ablation Studies
Sample Efficiency of Ranked Choice Training Method. We have evaluated the generalization performance as the amount of training data increases. For reference, when using all training data, the LC and WR for DPO are 41.24 and 40.24, respectively. As shown in Figure 3, Mallows-RMJ-PO-Top-2 can achieve comparable performance to full-data DPO when trained on much smaller data. Notably, strong performance can already be observed when using roughly 40% of the full dataset. This shows that our ranked choice method can achieve high sample efficiency.
Figure 3: AlpacaEval 2 metrics evolution with training steps for Mallows-RMJ-PO-Top-2.
Impact of $\beta$ on generalization performance. Figure 4 shows the $\beta$ sensitivity.
Figure 4: Effect of tuning $\beta$ for Mallows-RMJ-PO-Top-2.
Impact of different treatments for ranked responses. Given a set of ranked responses, one can (i) use only the top–bottom pair (Meng et al., 2024; Chen et al., 2025; Zhao et al., 2024; Gupta et al., 2025), (ii) decompose the ranking into all possible pairwise comparisons (Ouyang et al., 2022; Zhang et al., 2024), or (iii) directly train on ranked choice data (ours). We analyze how these different treatments affect both generalization performance and training efficiency.
As shown in Table 10, our methods achieve the best generalization performance with only a modest increase in training time. In contrast, decomposing the ranking into pairs greatly inflates the data size, resulting in much longer training time while yielding the worst performance.
Table 10: Evaluation Results of Different Treatments for Ranked Responses.
Method AlpacaEval 2 LC (%) AlpacaEval 2 WR (%) Arena-Hard WR (%) Training Time (Hours)
DPO 41.24 (0.35) 40.24 (1.66) 32.6 (30.6, 34.7) 1.5
DPO-AllPairs 33.02 (0.24) 38.47 (1.65) 29.6 (27.3, 31.5) 11.5
MNL-PO-Top-2 40.12 (0.22) 47.69 (1.67) 35.8 (33.2, 38.2) 3.5
Mallows-RMJ-PO-Top-2 41.95 (0.26) 53.01 (1.68) 37.2 (35.0, 39.4) 3.5
  • Notes: DPO and DPO-AllPairs correspond to treatment (i) and (ii), respectively. Standard errors (in parentheses) follow AlpacaEval 2 metrics, and 95% confidence intervals follow Arena-Hard WR.
@@ Appendix G Qualitative Examples
  • Notes: In the Excel test, Mallows-RMJ-PO-Top-2 answered all three questions correctly, whereas DPO incorrectly calculated the median in the last question using the average function.
@@
Table 14: Qualitative Example (Factual Explanatory Question)
Question: Do you know why turkeys became the official food of thanksgiving?
Mallows-RMJ-PO-Top-2 DPO
The origins of turkey as the centerpiece of the modern American Thanksgiving meal are not entirely clear, but there are several theories. Here’s a brief history: The Wampanoag and the Pilgrims: The first Thanksgiving, which took place in 1621, was a harvest feast celebrated by the Pilgrims, early European settlers of Plymouth Colony, and the Wampanoag Native American tribe. The Wampanoag had helped the Pilgrims adapt to the new land and taught them how to farm. The feast likely featured foods native to the region, such as venison, corn, squash, and beans. There is no direct evidence that turkey was served at this specific feast. Wild Turkeys in North America: Wild turkeys (Meleagris gallopavo) were abundant in North America, and Native Americans had hunted them for centuries. The Pilgrims would have been familiar with wild turkeys, which were a common food source in the region. 17th-Century European Traditions: In Europe, especially in England, roasted capon (a castrated male chicken) or goose was a traditional centerpiece for special occasions like harvest festivals and celebrations. When European settlers arrived in North America, they brought these culinary traditions with them. 18th and 19th Centuries: The Rise of Turkey as a Symbol of Thanksgiving: Over time, turkey became a popular choice for special occasions in America, including Thanksgiving. Several factors contributed to its rise: 1. Availability: Domesticated turkeys were introduced to North America in the early 1600s and became widely available in the 18th century. 2. Marketing and Promotion: In the mid-1800s, turkey farmers and butchers began promoting turkey as a centerpiece for Thanksgiving, capitalizing on its association with the holiday. 3. Sarah Josepha Hale’s Influence: Sarah Josepha Hale, a prominent American writer and editor, campaigned for a national Thanksgiving holiday through her magazine, Godey’s Lady’s Book. In 1827, she published a recipe for roasted turkey with stuffing, which helped popularize the dish. 4. Abraham Lincoln’s Proclamation: In 1863, during the Civil War, President Abraham Lincoln declared Thanksgiving a national holiday, and turkey became an integral part of the celebration. His wife, Mary Todd Lincoln, was known to serve roasted turkey at the White House. By the early 20th century, turkey had become the traditional main course for Thanksgiving in the United States, largely due to a combination of these factors. While it’s unlikely that turkey was the primary protein at the first Thanksgiving, it has since become an integral part of the holiday’s culinary tradition. A common misconception! Turkeys didn’t actually become the official food of Thanksgiving because of some historical decree. Instead, wild turkeys were a natural part of the traditional diet and celebrations of early American settlers, particularly the Pilgrims and Wampanoag Native Americans. Here’s the backstory: In 1621, the Pilgrims, who had established the Plymouth Colony, held a harvest feast to celebrate their first successful crop yield. They invited the Wampanoag tribe, who had helped them adapt to the new land and taught them agricultural techniques. The feast, now known as the First Thanksgiving, lasted for three days and featured a variety of foods, including: 1. Venison (deer meat) 2. Wild fowl (likely including ducks, geese, and passenger pigeons) 3. corn (maize) 4. Squash 5. Beans 6. Seafood (like cod and eel) Turkeys were present, but not the centerpiece. Wild turkeys were abundant in the region, and they would have been a natural addition to the harvest meal, along with other game meats. Over time, as European customs merged with American traditions, roasted turkey became a staple of annual harvest celebrations, including Thanksgiving, which was not officially declared a national holiday until 1863 by President Abraham Lincoln. By then, turkey had already become a popular choice due to its availability, affordability, and symbolic association with abundance and gratitude. So, while turkeys weren’t specifically designated as the “official food” of Thanksgiving, their presence at the First Thanksgiving and their subsequent popularity in American cuisine cemented their place as a classic Thanksgiving dish.
  • Notes: These outputs are illustrative rather than fact-checked. Mallows-RMJ-PO-Top-2 responds by providing a historical narrative aligned with the question’s framing, whereas DPO places greater emphasis on questioning and qualifying the premise of the question.
Table 15: Qualitative Example (Simplification)
Question: Simplify “Most of the basic functions of current virtualization software are inherited from this early VM OS. The virtualization software is now represented with the term hypervisor. Hardware virtualization is used to share the resources of the cloud providers effectively with the customers by generating an illusion of dedicated computing, storage, and networking on a computing infrastructure. The concept of virtualization is physically implemented using the hypervisor modules, and the operation and processing of hypervisors are materialized by the OSS. In other words, hypervisor modules are installed on top of OSS, which act as an interface between hardware units and hypervisor packages.”
Mallows-RMJ-PO-Top-2 DPO
Here’s a simplified version: “Virtualization software, now called hypervisors, inherited many basic functions from early virtual machine operating systems. Hypervisors use hardware virtualization to provide customers with a dedicated computing, storage, and networking experience in the cloud, while sharing resources efficiently. The concept of virtualization is implemented through hypervisor modules, which are built on top of Operating System Software (OSS). OSS acts as an interface between hardware and hypervisor packages, enabling the operation and processing of hypervisors.” I removed some redundant phrases, rephrased sentences for better clarity, and condensed the text to preserve the main ideas. Let me know if you’d like any further simplification! Here’s a simplified version: Virtualization technology, now known as a hypervisor, originated from early VM OS fundamentals. It enables cloud providers to efficiently share resources with customers by creating a simulated dedicated computing, storage, and networking experience. Hypervisor modules, built on Open-Source Software (OSS), act as an interface between hardware and hypervisor packages, physically implementing virtualization concepts.
  • Notes: These outputs are illustrative rather than fact-checked. Mallows-RMJ-PO-Top-2 exhibits higher fidelity to the source text, whereas DPO introduces semantic distortion during simplification.